OpenStudy (anonymous):

Choose the value of the x determinant (Dx) in the system. 3x – 7y = 11 2x + 5y = –4 Answer A. 29 B. 83 C. 1 D. 27 Choose the value of the x determinant (Dx) in the system. 3x – 7y = 11 2x + 5y = –4 Answer A. 29 B. 83 C. 1 D. 27 @Mathematics

OpenStudy (anonymous):

What is the value for the coefficient determinant (D) in the following system? 2x + y = 2 3x + 4y = –7 For the coefficient determinant, make the coefficients of x one column and the coefficients of y the second. First Row (2, 1) Second Row(3, 4) To evaluate the determinant, use this formula for determinant with Row 1: (a, b) Row 2 (c, d) Det = ad - bc So for this problem = 2*4 - 3*1 = 5

OpenStudy (anonymous):

opps wrong question :P

OpenStudy (amistre64):

Dx I believe is the numerator of the Cramer Rule. Such that Dx is: -4(-7) -11(5)